Mrs. Simpson buys loaves of bread and quarts of milkhh each week at prices of S and 80 cents, respectively. At present she is buying these products in amounts such that the marginal utilities from the last units purchased of the two products are 80 and 70 utils, respectively. Is she buying the utility-maximizing combination of bread and milk? If not, how should she reallocate her expenditures between the two goods?
Answer #1

Marginal rate of transformation (MRT) should be equal to Marginal rate of substitution (MRS) for the consumer to be in equilibrium.

MRT=P1/P2=1/.8=1.25

MRS=MU1/MU2=80/70=1.14

Since MRT>MRS she is not buying the right amount of goods.

MU1/MU2

MU1/P1

We know from the law of diminishing marginal utility that if the consumption of a good rises,the marginal utility from it falls.

If the consumption of good 2 is increased,MU2 will fall and eventually MU2/P2 will be equal to MU1/P1

Thus,she should buy more units of good 2 or milk.